# Mechanical quadrature squeezing beyond the 3dB limit via quantum learning control

### v1

@@ Abstract
The preparation of mechanical quadrature-squeezed states holds significant importance in cavity optomechanics because the squeezed states have extensive applications in elucidating fundamental quantum mechanics and exploiting modern quantum technonogy. Here, we propose a reliable scheme for generating mechanical quadrature squeezing in a typical cavity optomechanical system via seeking for optimal cavity-field driving pulses using a quantum-learning-control method. We realize strong quadrature squeezing exceeded the 3 dB steady-state limit in the mechanical resonator subjected to thermal noise with one hundred thermal phonons. Furthermore, the mechanical squeezing can be ultrafastly created within one mechanical oscillation period. We also obtain the optimal pulsed drivings associated with the created mechanical squeezings and analyze the mechanism for mechanical squeezing generation. This work will motivate succeeding applications of optimal quantum control in quantum optics and quantum information science.
: New J. Phys.
Keywords: optomechanical systems, pulsed driving, mechanical quadrature squeezing, gradient descent algorithm, quantum learning control
1 Introduction
@@
Cavity optomechanics [1] is a frontier research field focusing on the radiation-pressure interaction between cavity fields and moving mirrors at macroscopic scales. Owing to the nonlinear properties of the optomechanical interaction and the development of laser cooling techniques [2, 3, 4, 5], optomechanical systems provide new paths and opportunities for generation and manipulation of macroscopic nonclassical states, such as entangled states [6, 7, 8, 9, 10, 11], quantum superposition states [12, 13, 14, 15], and squeezed states of fields [16, 17, 18, 19, 20, 21] and mechanical resonators [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43].
@@
Quantum squeezing, as an essential quantum resource, plays an increasingly significant role in modern quantum technology ranging from the detection of weak forces and fields to quantum information processing [44, 45]. Quantum squeezing of mechanical modes holds significance as it has the potential to enhance the accuracy of quantum measurements [46]. There exist many theoretical and experimental schemes for generating mechanical squeezing based on various methods, such as parametric squeezing [24, 25, 26, 27, 28, 29, 30, 31], quantum measurement [32, 33, 34, 35, 36], quantum state transfer [37, 38], mechanical nonlinearities [24, 39, 40], and reservoir engineering [42, 43]. Although great advances have been made in squeezing generation, how to generate mechanical squeezing beyond the 3 dB limit [38, 26, 47] in a typical optomechanical system is still an important and desired task in this field. In addition, from the view point of realistic application, an ultrafast generation of strong mechanical squeezing remains a challenge.
@@
In cavity optomechanical systems, the external optical driving provides an effective way to control the quantum properties and dynamical behaviors of the system [48, 49, 25, 27, 50]. The design of an optimal optical driving to achieve a supposed goal is an interesting topic in this system. Currently, quantum control techniques have been successfully applied to various schemes in cavity optomechanical systems [51], such as robust state transfer [52], optimized cooling [53, 54, 55], entanglement enhancement [56, 57], and strong squeezing [58, 59]. Furthermore, quantum learning control [60] has been recognized as a powerful way to accomplish the complex control tasks, and it is expected to provide new ways to optimize the control of cavity optomechanical systems.
@@
In this paper, we apply the quantum-learning-control method to prepare mechanical quadrature squeezing in a typical cavity optomechanical system. Our scheme can break the 3 dB steady-state limit even when the environment thermal phonon occupation associated with the mechanical resonator reaches to the order of one hundred. The mechanical quadrature squeezing is achieved by designing a proper pulsed driving to the cavity field. Concretely, we use the gradient descent algorithm to iteratively optimize the pulsed driving for achieving strong mechanical squeezing. The optimal waveforms of the pulsed driving amplitude and phase are also obtained. In particular, mechanical squeezing can be ultrafastly prepared within one mechanical oscillation period, providing more chance for realistic applications of the generated squeezing before decoherence. Our method will motivate succeeding works on optimal quantum control in cavity optomechanical systems.
@@
The rest of this paper is organized as follows. In Sec. 2, we introduce the Hamiltonians and present the equations of motion. In Sec. 3, we show the generation of mechanical quadrature squeezing using quantum-learning control method. In Sec. 4, we present some discussions on the physical mechanism and experimental implementation of this scheme and conclude this work. An Appendix is presented to show the derivation of the variation $\delta[\Delta X_{b}^{2}(\theta,T)]/\delta\mathcal{Q}_{m}$ in gradient descent algorithm.
2 Hamiltonians and equations of motion
@@
$H_{\mathrm{opt}}(t)$ $=\Delta_{\mathrm{c}}a^{{\dagger}}a+\omega_{\mathrm{m}}b^{{\dagger}}b-g_{0}a^{{\dagger}}a(b^{{\dagger}}+b)+[\Omega(t)\mathrm{e}^{-i\phi(t)}a^{{\dagger}}+\mathrm{H.c.}],$ (1)
where $a^{{\dagger}}$ $(a)$ and $b^{{\dagger}}$ $(b)$ are, respectively, the creation (annihilation) operators of the cavity field and the mechanical resonator, with the corresponding resonance frequencies $\omega_{\mathrm{c}}$ and $\omega_{\mathrm{m}}$. The $\Delta_{\mathrm{c}}=\omega_{\mathrm{c}}-\omega_{\mathrm{L}}$ is the detuning of the cavity-field resonance frequency with respect to the pulsed-field carrier frequency. The $g_{0}$ term describes the radiation-pressure coupling between the cavity field and the mechanical resonator, with $g_{0}$ being the single-photon optomechanical coupling strength.
@@
where $\rho$ is the density matrix of the optomechanical system, $\mathcal{D}[o]\rho=o\rho o^{{\dagger}}-(o^{{\dagger}}o\rho+\rho o^{{\dagger}}o)/2$ (for $o=a$, $a^{{\dagger}}$, $b$, and $b^{{\dagger}}$) is the Lindblad superoperator [61], and $H_{\mathrm{opt}}(t)$ is defined in Eq. (1). The parameters $\kappa$ and $\gamma$ are, respectively, the damping rates of the cavity field and the mechanical oscillation, and $\bar{n}_{\mathrm{m}}$ is the environmental thermal-excitation occupation of the mechanical resonator.
Considering the strong-driving case of the optomechanical cavity, then the dynamics of the system can be linearized. To this end, we adopt the displacement-transformation method to separate the semiclassical motion and the quantum fluctuation. Concretely, we perform the displacement transformations $D_{a}(\alpha)=\exp(\alpha a^{\dagger}-\alpha^{\ast}a)$ and $D_{b}(\beta)=\exp(\beta b^{\dagger}-\beta^{\ast}b)$ for the density operator $\rho(t)$, namely,
$\rho^{\prime}(t)=D_{a}(\alpha)D_{b}(\beta)\rho(t)D_{b}^{\dagger}(\beta)D_{a}^{\dagger}(\alpha).$ (3)
By substituting Eq. (3) into Eq. (2) and assuming that the coefficients of the driving terms are zero, we obtain the quantum master equation in the displaced representation as
@@
$H_{\mathrm{dis}}(t)$ $=\Delta(t)a^{\dagger}a+\omega_{\mathrm{m}}b^{\dagger}b+[G(t)a^{\dagger}+G^{\ast}(t)a](b+b^{\dagger})-g_{0}a^{\dagger}a(b+b^{\dagger}).$ (5)
In the strong-driving case, $|\alpha(t)|\gg 1$, we can safely omit the $g_{0}$ term in Eq. (5), and obtain the linearized Hamiltonian
@@
where we introduce the linearized optomechanical-coupling strength $G(t)=g_{0}\alpha(t)$ and the normalized driving detuning $\Delta(t)=\Delta_{c}+g_{0}[\beta(t)+\beta^{*}(t)]$. Here, the two displacement amplitudes $\alpha(t)$ and $\beta(t)$ are governed by the equations of motion,
$\dot{\alpha}(t)$ $=-[i\Delta(t)+\kappa/2]\alpha(t)+i\Omega(t)\text{e}^{-i\phi(t)},$ (7a)
$\dot{\beta}(t)$ $=-(i\omega_{\mathrm{m}}+\gamma/2)\beta(t)-ig_{0}|\alpha(t)|^{2}.$ (7b)
@@
The dynamic properties of the linearized optomechanical system are completely described by both the first- and second-order moments of the system operators. Using the relation $\partial_{t}\langle o_{i}o_{j}\rangle=\text{Tr}(\dot{\rho}^{\prime}o_{i}o_{j})$ for $o_{i},o_{j}$in {$a$, $a^{\dagger}$, $b$, $b^{\dagger}$} and Eq. (4) under the replacement of $H_{\mathrm{dis}}(t)\rightarrow H_{\mathrm{lin}}(t)$, we obtain the equations of motion of all these second-order moments, which can be expressed as a compact form as
@@
where $\mathbf{X}(t)=[\langle a^{\dagger}a\rangle,\langle b^{\dagger}b\rangle,\langle a^{\dagger}b\rangle,\langle ab^{\dagger}\rangle,\langle a^{\dagger}a^{\dagger}\rangle,\langle a^{\dagger}b^{\dagger}\rangle,$ $\langle b^{\dagger}b^{\dagger}\rangle,\langle aa\rangle,\langle ab\rangle,\langle bb\rangle]^{\text{T}}$ (“$\mathrm{T}$” denotes the matrix transpose), $\mathbf{N}(t)=[0,\gamma\bar{n}_{\mathrm{m}},0,0,0,ig_{0}\alpha^{\ast},0,0,-ig_{0}\alpha,0]^{\mathrm{T}}$, and the coefficient matrix is introduced as $\mathbf{M}(t)=\left(\begin{array}[]{cc}\mathbf{H}&\mathbf{I}\\ \mathbf{J}&\mathbf{K}\end{array}\right)$, with
@@
$K_{1}$ $=i\Delta(t)-i\omega_{\mathrm{m}}-(\kappa+\gamma)/2,$ (10a)
@@
$K_{3}$ $=i\Delta(t)+i\omega_{\mathrm{m}}-(\kappa+\gamma)/2,$ (10c)
@@
3 Generation of mechanical quadrature squeezing
where $\mathbf{u}(t)=[X_{a}(t),Y_{a}(t),X_{b}(t),Y_{b}(t)]^{\mathrm{T}}$. In the present case, the expectation values $\langle\mathbf{u}_{i}(t)\rangle$ and $\langle\mathbf{u}_{j}(t)\rangle$ are zero and $\langle\mathbf{u}_{i}(t)\mathbf{u}_{j}(t)\rangle$ is a linear function of these second-order moments. To describe the quadrature squeezing, we further introduce the rotating-quadrature operator $X_{b}(\theta,t)\equiv X_{b}(t)\cos\theta+Y_{b}(t)\sin\theta$ ($\theta$ is the squeezing angle). The variance of the rotating-quadrature operator is given by
$\Delta X_{b}^{2}(\theta,t)$ $=\mathbf{V}_{33}(t)\cos^{2}\theta+\mathbf{V}_{44}(t)\sin^{2}\theta+\frac{1}{2}[\mathbf{V}_{34}(t)+\mathbf{V}_{43}(t)]\sin(2\theta).$ (12)
@@ 3 Generation of mechanical quadrature squeezing
   1. Perform the gradient descent algorithm $\mathcal{Q}_{m+1}\!=\!\mathcal{Q}_{m}-\chi_{\mathcal{Q}}\{\delta[\Delta X_{b}^{2}(\theta,T)]/\delta\mathcal{Q}_{m}$} ($\mathcal{Q}$ represents either $\Omega$ or $\phi$, $\chi_{\mathcal{Q}}$ is the learning rate, and $m$ indicates the iteration number) to iteratively minimize $\Delta X_{b}^{2}(\theta,T)$.
@@
The mechanical quadrature squeezing can be well quantized by the squeezing degree [38]

### v2

# Optimized mechanical quadrature squeezing beyond the 3 dB limit via gradient-descent algorithm

@@ Abstract
The preparation of mechanical quadrature-squeezed states holds significant importance in cavity optomechanics because the squeezed states have extensive applications in understanding fundamental quantum mechanics and exploiting modern quantum technonogy. Here, we propose a reliable scheme for generating mechanical quadrature squeezing in a typical cavity optomechanical system via seeking for optimal cavity-field driving pulses using the gradient-descent algorithm. We realize strong quadrature squeezing in a mechanical resonator that exceeds the $3$ dB steady-state limit, even with a thermal phonon occupancy of one hundred. Furthermore, the mechanical squeezing can be ultrafastly created within one mechanical oscillation period. We also obtain the optimal pulsed drivings associated with the created mechanical squeezings and analyze the mechanism for mechanical squeezing generation. This work will promote the application of optimal quantum control in quantum optics and quantum information science.
I INTRODUCTION
@@
Cavity optomechanics Kippenberg2014RMP is a frontier research field focusing on the radiation-pressure interaction between cavity fields and moving mirrors at macroscopic scales. Owing to the nonlinear properties of the optomechanical interaction and the development of laser cooling techniques Wilson-Rae2007PRL ; Marquardt2007PRL ; ChanNature2011 ; TeufelNature2011 , optomechanical systems provide new paths and opportunities for generation and manipulation of macroscopic nonclassical states, such as entangled states VitaliPRL2007 ; Palomaki2013Science ; Riedinger2016Nature ; Jiao2020PRL ; Yu2020Nature ; Lai2022PRL , quantum superposed states Bose1997PRA ; Marshall2003PRL ; LiaoPRL2016 ; LiaoPRA2020 , and squeezed states of fields FabrePRA1994 ; ManciniPRA1994 ; Brooks2012 ; PurdyPRx2013 ; Safavi-Naeini2013 ; AgarwalNJP2014 and mechanical resonators WoolleyPRA2008 ; HuangPRA2023 ; TeufelNN2009 ; NunnenkampPRA2010 ; LiaoPRA2011 ; MilburnOL1981 ; MariPRL2009 ; PontinPRL2014 ; PontinPRL2016 ; ChowdhuryPRL2020 ; YongChun2023PRA ; ClerkNJP2008 ; SzorkovszkyPRL2011 ; VasilakisNP2015 ; MengPRL2020 ; PirkkalainenPRL2015 ; GenesPRA2009 ; AgarwalPRA2016 ; AsjadPRA2014 ; LuPRA2015 ; RashidPRL2016 ; KronwaldPRA2013 ; XiongOL2018 .
@@
Quantum squeezing, as an essential quantum resource, plays an increasingly significant role in modern quantum technology ranging from quantum precision measurement to quantum information processing Braunstein2003 ; Drummond2004 . Quantum squeezing of mechanical modes holds significance as it has the potential to enhance the accuracy of quantum measurements CavesRMP1980 . There exist many theoretical and experimental schemes for generating mechanical squeezing based on various methods, such as parametric squeezing NunnenkampPRA2010 ; LiaoPRA2011 ; MilburnOL1981 ; MariPRL2009 ; PontinPRL2014 ; PontinPRL2016 ; ChowdhuryPRL2020 ; YongChun2023PRA , quantum measurement ClerkNJP2008 ; SzorkovszkyPRL2011 ; VasilakisNP2015 ; MengPRL2020 ; PirkkalainenPRL2015 , quantum state transfer GenesPRA2009 ; AgarwalPRA2016 ; HuangPRA2023 , mechanical nonlinearities NunnenkampPRA2010 ; AsjadPRA2014 ; LuPRA2015 , and reservoir engineering KronwaldPRA2013 ; XiongOL2018 . Although great advances have been made in squeezing generation, how to generate mechanical squeezing beyond the 3 dB limit AgarwalPRA2016 ; MilburnOL1981 ; Scully1997 in a typical optomechanical system is still an important and desired task in this field. In addition, from the view point of realistic application, an ultrafast generation of strong mechanical squeezing remains a challenge.
@@
In cavity optomechanical systems, the external optical driving provides an effective way to control the quantum properties and dynamical behaviors of the system VannerPNAS2011 ; VannerNC2013 ; LiaoPRA2011 ; MariPRL2009 ; MeenehanPRX2015 . The design of an optimal optical driving to achieve a supposed goal is an interesting topic in this system. Currently, quantum control techniques have been successfully applied to various schemes in cavity optomechanical systems Hammerer2017AIMO , such as robust state transfer Wang2021AP , optimized cooling Jacobs2011 ; MagriniNature2021 ; LiuPRA2022 , entanglement enhancement Stefanatos2017QST ; ClarkeNJP2020 , and strong squeezing Xiong2020PR ; Basilewitsch2019AQE . Furthermore, optimal control technology based on the gradient-descent algorithm Dong2020 has been recognized as a powerful method to accomplish complex control tasks, and hence it is expected to provide new ways to optimize the control of cavity optomechanical systems.
@@
In this paper, we apply the gradient-descent algorithm to prepare mechanical quadrature squeezing in a typical cavity optomechanical system. Our scheme can break the $3$ dB steady-state squeezing limit even when the environment thermal phonon occupation associated with the mechanical resonator reaches the order of one hundred. The mechanical quadrature squeezing is achieved by designing a proper pulsed driving to the cavity field. Concretely, we use the gradient-descent algorithm to iteratively optimize the pulsed driving for achieving strong mechanical squeezing. The optimal waveforms of the pulsed driving amplitude and phase are also obtained. In particular, mechanical squeezing can be ultrafastly prepared within one mechanical oscillation period, providing more chance for realistic applications of the generated squeezing before decoherence. Our method will encourage further researchs on optimal quantum control in cavity optomechanical systems.
@@
The rest of this paper is organized as follows. In Sec. II, we introduce the Hamiltonians and present the equations of motion. In Sec. III, we show the generation of mechanical quadrature squeezing using the gradient-descent algorithm. In Sec. IV, we present some discussions on the physical mechanism and experimental implementation of this scheme and conclude this work. An Appendix is presented to show the derivation of the variation $\delta[\Delta X_{b}^{2}(\theta,T)]/\delta\mathcal{Q}_{m}$ used in the gradient-descent algorithm.
II HAMILTONIANS AND EQUATIONS OF MOTION
@@
$H_{\mathrm{opt}}(t)$ $=\Delta_{\mathrm{c}}a^{{\dagger}}a+\omega_{\mathrm{m}}b^{{\dagger}}b-g_{0}a^{{\dagger}}a(b^{{\dagger}}+b)$
$\hskip 8.5359pt+[\Omega(t)\mathrm{e}^{-i\phi(t)}a^{{\dagger}}+\mathrm{H.c.}],$ (1)
where $a^{{\dagger}}$ $(a)$ and $b^{{\dagger}}$ $(b)$ are, respectively, the creation (annihilation) operators of the cavity field and the mechanical resonator, with the corresponding resonance frequencies $\omega_{\mathrm{c}}$ and $\omega_{\mathrm{m}}$. The $\Delta_{\mathrm{c}}=\omega_{\mathrm{c}}-\omega_{\mathrm{L}}$ is the detuning of the cavity-field resonance frequency with respect to the pulsed-field carrier frequency. The $g_{0}$ term describes the radiation-pressure coupling between the cavity field and the mechanical resonator, with $g_{0}$ being the single-photon optomechanical-coupling strength.
@@
where $\rho$ is the density matrix of the optomechanical system, $\mathcal{D}[o]\rho=o\rho o^{{\dagger}}-(o^{{\dagger}}o\rho+\rho o^{{\dagger}}o)/2$ (for $o=a$, $a^{{\dagger}}$, $b$, and $b^{{\dagger}}$) is the Lindblad superoperator Agarwal2013 , and $H_{\mathrm{opt}}(t)$ is defined in Eq. (II). The parameters $\kappa$ and $\gamma$ are, respectively, the damping rates of the cavity field and the mechanical oscillation, and $\bar{n}_{\mathrm{m}}$ is the environmental thermal-excitation occupation of the mechanical resonator.
Considering the strong-driving case of the optomechanical cavity, then the dynamics of the system can be linearized. To this end, we adopt the displacement-transformation method to separate the semiclassical motion and quantum fluctuation. Concretely, we perform the displacement transformations $D_{a}(\alpha)=\exp(\alpha a^{\dagger}-\alpha^{\ast}a)$ and $D_{b}(\beta)=\exp(\beta b^{\dagger}-\beta^{\ast}b)$ for the density operator $\rho(t)$, namely,
$\rho^{\prime}(t)=D_{a}(\alpha)D_{b}(\beta)\rho(t)D_{b}^{\dagger}(\beta)D_{a}^{\dagger}(\alpha),$ (3)
where $\rho^{\prime}(t)$ represents the density operator in the displaced representation, $\alpha(t)$ and $\beta(t)$ are the displacement amplitudes of the cavity field and the mechanical resonator, respectively. By substituting $\rho(t)=D_{b}^{\dagger}(\beta)D_{a}^{\dagger}(\alpha)\rho^{\prime}(t)D_{a}(\alpha)D_{b}(\beta)$ into Eq. (2) and setting the coefficients of the driving terms to be zero, we obtain the quantum master equation in the displaced representation as
@@
$H_{\mathrm{dis}}(t)$ $=\Delta(t)a^{\dagger}a+\omega_{\mathrm{m}}b^{\dagger}b-g_{0}a^{\dagger}a(b+b^{\dagger})$
$\hskip 8.5359pt+[G(t)a^{\dagger}+G^{\ast}(t)a](b+b^{\dagger}).$ (5)
In Eq. (5), we introduce the linearized optomechanical-coupling strength $G(t)=g_{0}\alpha(t)$ and the normalized driving detuning $\Delta(t)=\Delta_{c}+g_{0}[\beta(t)+\beta^{*}(t)]$. The two displacement amplitudes $\alpha(t)$ and $\beta(t)$ are governed by the equations of motion,
$\dot{\alpha}(t)$ $=-\left[i\Delta(t)+\frac{\kappa}{2}\right]\alpha(t)+i\Omega(t)\text{e}^{-i\phi(t)},$ (6a)
$\dot{\beta}(t)$ $=-\left(i\omega_{\mathrm{m}}+\frac{\gamma}{2}\right)\beta(t)-ig_{0}|\alpha(t)|^{2}.$ (6b)
In the strong-driving case, $|\alpha(t)|\gg 1$, we can safely omit the $g_{0}$ term in Eq. (5), and obtain the linearized optomechanical Hamiltonian
@@
The dynamic properties of the linearized optomechanical system are completely described by both the first- and second-order moments of the system operators. Using the relation $\partial_{t}\langle o_{i}o_{j}\rangle=\text{Tr}(\dot{\rho}^{\prime}o_{i}o_{j})$ for $o_{i},o_{j}$in {$a$, $a^{\dagger}$, $b$, $b^{\dagger}$} and Eq. (4) under the replacement of $H_{\mathrm{dis}}(t)\rightarrow H_{\mathrm{lin}}(t)$, we obtain the equations of motion of all these second-order moments, which can be expressed as
@@
where $\mathbf{X}(t)=(\langle a^{\dagger}a\rangle,\langle b^{\dagger}b\rangle,\langle a^{\dagger}b\rangle,\langle ab^{\dagger}\rangle,\langle a^{\dagger}a^{\dagger}\rangle,\langle a^{\dagger}b^{\dagger}\rangle,$ $\langle b^{\dagger}b^{\dagger}\rangle,\langle aa\rangle,\langle ab\rangle,\langle bb\rangle)^{\text{T}}$ (“$\mathrm{T}$” denotes the matrix transpose), $\mathbf{N}(t)=(0,\gamma\bar{n}_{\mathrm{m}},0,0,0,ig_{0}\alpha^{\ast},0,0,-ig_{0}\alpha,0)^{\mathrm{T}}$, and the coefficient matrix is introduced as $\mathbf{M}(t)=\left(\begin{array}[]{cc}\mathbf{H}&\mathbf{I}\\ \mathbf{J}&\mathbf{K}\end{array}\right)$, with
@@
$K_{1}$ $=i\Delta(t)-i\omega_{\mathrm{m}}-\frac{1}{2}(\kappa+\gamma),$ (10a)
@@
$K_{3}$ $=i\Delta(t)+i\omega_{\mathrm{m}}-\frac{1}{2}(\kappa+\gamma)\,$ (10c)
@@
   1. Perform the gradient-descent algorithm $\mathcal{Q}_{m+1}\!=\!\mathcal{Q}_{m}-\chi_{\mathcal{Q}}\{\delta[\Delta X_{b}^{2}(\theta,T)]/\delta\mathcal{Q}_{m}$} ($\mathcal{Q}$ represents either $\Omega$ or $\phi$, $\chi_{\mathcal{Q}}$ is the learning rate, and $m$ indicates the iteration number) to iteratively minimize $\Delta X_{b}^{2}(\theta,T)$.
@@
III GENERATION OF MECHANICAL QUADRATURE SQUEEZING
where $\mathbf{u}(t)=(X_{a}(t),Y_{a}(t),X_{b}(t),Y_{b}(t))^{\mathrm{T}}$. In the present case, the expectation values $\langle\mathbf{u}_{i}(t)\rangle$ and $\langle\mathbf{u}_{j}(t)\rangle$ are zero and $\langle\mathbf{u}_{i}(t)\mathbf{u}_{j}(t)\rangle$ is a linear function of these second-order moments. To describe the quadrature squeezing, we further introduce the rotating-quadrature operator $X_{b}(\theta,t)\equiv X_{b}(t)\cos\theta+Y_{b}(t)\sin\theta$ ($\theta$ is the squeezing angle). The variance of the rotating-quadrature operator is given by
$\Delta X_{b}^{2}(\theta,t)$ $=\mathbf{V}_{33}(t)\cos^{2}\theta+\mathbf{V}_{44}(t)\sin^{2}\theta$ (12)
$+\frac{1}{2}[\mathbf{V}_{34}(t)+\mathbf{V}_{43}(t)]\sin(2\theta).$
The mechanical quadrature squeezing can be well quantized by the squeezing degree AgarwalPRA2016